OpenStudy (anonymous):

What is this equation in simplified radical form? the square root of 98xw^8

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow \sqrt{98xw^8} = w^4\sqrt{7*7*2*x} = 7w^4\sqrt{2x} }\)

OpenStudy (anonymous):

thanks